OpenStudy (anonymous):

Let f(x) = x2 – 81. Find f–1(x).

OpenStudy (anonymous):

f(x)=x^2-81 let's define f(x)=y

OpenStudy (anonymous):

y=x^2-81

OpenStudy (anonymous):

now switch variable

OpenStudy (anonymous):

x=y^2-81 solve for y

OpenStudy (anonymous):

I'm not sure how too

OpenStudy (anonymous):

start by addng 81 to both sides

OpenStudy (anonymous):

x + 81= y2

OpenStudy (anonymous):

??

OpenStudy (anonymous):

good

OpenStudy (anonymous):

y^2=x+81 now take square root of both sides

OpenStudy (anonymous):

+- x + 81?

OpenStudy (anonymous):

I think you meant \[\pm\sqrt{x+81}\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

yep, that's it; that's f^-1(x)

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

you are welcome